Differentiate the following function
g(x) = e
−5x +

1
x
− tan x
3 p
x
2 + 1 
.
1
Expert's answer
2020-12-13T17:12:28-0500
"g'(x)=\bigg(e^{-5x}+(\dfrac{1}{x}-\tan x)^3\sqrt{x^2+1}\bigg)'"

"=-5e^{-5x}+3(\dfrac{1}{x}-\tan x)^2(-\dfrac{1}{x^2}-\dfrac{1}{\cos^2x})\sqrt{x^2+1}"

"+(\dfrac{1}{x}-\tan x)^3\dfrac{x}{\sqrt{x^2+1}}"